OpenStudy (anonymous):

Simplify (ab)(ab)(ab)(ab) ab^4?

jimthompson5910 (jim_thompson5910):

it would be (ab)^4 but remember the 4 applies to both a and b, so it would be a^4b^4

jimthompson5910 (jim_thompson5910):

ie there are 4 copies of b being multiplied, so b*b*b*b = b^4 and there are 4 copies of a being multiplied, so a*a*a*a = a^4

OpenStudy (anonymous):

yes that is correct

jimthompson5910 (jim_thompson5910):

put together \[\large (ab)(ab)(ab)(ab) = (ab)^4 = a^4b^4\]

OpenStudy (anonymous):

yup thats right

OpenStudy (anonymous):

Thank you, I see what you're explaining.

jimthompson5910 (jim_thompson5910):

you're welcome